# Dual membership certificatesJoonyeob Lee, Dávid Papp, Anita Varga

### v2

@@ 7 Application to polynomial optimization
In the presence of strong duality, if we are only concerned about the approximate numerical value of this lower bound, we may opt to solve the dual problem of (29) instead; for example, if $\mathcal{C}$ is a suitable SOS cone, (29) is known as the SOS relaxation of the POP (27), and the dual of (29) is known as Lasserre’s moment relaxation [20].
The theoretical conundrum is now similar to the issue we have encountered in the Introduction and Section 6: every cone $\mathcal{C}$ that is typically used in this context to approximate $\mathscr{P}_{\Delta}$, including each of those examined later in this section, is represented as a linear image $\mathcal{K}_{A}$ of a “nice” cone $\mathcal{K}$ (recall Eq. (9)), which means that a typical numerical method solving (29) will result only in an approximate certificate, rigorously proving that some polynomial close to, but not identical to $p-\alpha$ is nonnegative over $\Delta$ [21]. On the dual side, a near-optimal feasible solution certifiably in $(\mathcal{C}^{*})^{\circ}$ can be found, but this only yields an upper bound on the optimal value of (29), that is, an upper bound on a lower bound on the optimal value of (27). In summary, both (29) and its dual may be solvable approximately using numerical methods, but neither numerical solutions provide a theoretically satisfactory, rigorously certified lower bound (with a nonnegativity certificate that is verifiable in rational arithmetic) on the optimal value of (27). Instead, additional post-processing, in the form of some rounding procedure, is required to turn the numerical bounds and certificates into rigorous ones, which (in addition to being additional work) puts additional constraints on the quality of the numerical solution. See, e.g., [31, 16, 21] for examples in the context of SOS certification.
@@ 7.3 Signomial functions and S-cones
Katthän, Naumann and Theobald have shown that, analogously to the SONC results, functions in $\mathcal{S}_{\mathscr{A},\mathscr{B}}$ can be written as sums of AG functions with support $(\mathscr{A},\mathscr{B})$ [17]. They have also characterized both even and odd AG functions in terms of at most two relative entropy cone inequalities [17, Theorem 2.8]. This representation is compatible with the concept of dual certificates. For example, an $n$-variate odd AG function given by (31) with support $(\mathscr{A},\{{\beta}\})$ is nonnegative on $\mathbb{R}^{n}$ if and only if there exists a $\nu\in\mathbb{R}_{+}^{\mathscr{A}}$ such that
@@ 7.4 Nonnegative bases, LP- and SOCP-representable subcones of SOS polynomials
    Polynomials with nonnegative coefficients in a fixed basis chosen to be trivially nonnegative on $\Delta$;
@@ 8 Discussion
We presented a theoretical framework of cone membership certificates that allows us to interpret vectors in the dual cone as certificates of membership in the primal cone. The new certification scheme builds on the classic self-concordance theory of Nesterov and Nemirovski [25], but it goes beyond the well-known result that the gradient map associated with an SCB defines a bijection between the interior of a cone and the interior of its dual. The key difference is that $\mathcal{H}(b)$ and $\mathcal{B}(b)$ are full-dimensional cones, which means that we can use numerical methods to identify any of the certificates, and still establish rigorous certification of membership. Similarly, in the case of a low-dimensional projection $\mathcal{K}_{A}$ of the high-dimensional cone $\mathcal{K}$, an exact preimage $x\in\mathcal{K}$ of a vector $b\in\mathcal{K}_{A}$ can be computed from any (numerically computed) dual certificate. The same infrastructure also allowed us to compute exact primal feasible solutions of general convex conic optimization problems from suitable dual feasible solutions, with duality gaps that approach zero as the dual feasible solution approaches optimality.
We have also established that for many cones, some already existing optimization methods (specifically, feasible dual and primal-dual IPAs operating in appropriate neighborhoods of the central path), can be used to compute dual certificates. For certain cones (e.g., linear images of symmetric cones), the existing dual methods automatically compute dual certificates, and these can now be extended to compute exact primal feasible solutions that converge to a primal optimal solution, complementing the dual feasible iterates of the method. For the most general case of nonsymmetric conic optimization, the IPA recently proposed by the authors [27] can be applied.

### v3

@@ 7 Application to polynomial optimization
In the presence of strong duality, if we were only concerned about the approximate numerical value of this lower bound, we may opt to solve the dual problem of (29) instead; for example, if $\mathcal{C}$ is a suitable SOS cone, (29) is known as the SOS relaxation of the POP (27), and the dual of (29) is known as the moment relaxation [22]. The theoretical conundrum is now similar to the issue we have encountered in the Introduction and Section 6: every cone $\mathcal{C}$ that is typically used in this context to approximate $\mathscr{P}_{\Delta}$, including each of those examined later in this section, is represented as a linear image $\mathcal{K}_{A}$ of a “nice” cone $\mathcal{K}$ (recall Eq. (9)), which means that a typical numerical method solving (29) will result only in an approximate certificate, rigorously proving that some polynomial close to, but not identical to $p-\alpha$ is nonnegative over $\Delta$ [23]. On the dual side, a near-optimal feasible solution certifiably in $(\mathcal{C}^{*})^{\circ}$ can be found, but this only yields an upper bound on the optimal value of (29), that is, an upper bound on a lower bound on the optimal value of (27). In summary, both (29) and its dual may be solvable approximately using numerical methods, but neither numerical solutions provide a theoretically satisfactory, rigorously certified lower bound (with a nonnegativity certificate that is verifiable in rational arithmetic) on the optimal value of (27). Instead, additional post-processing, in the form of some rounding procedure, is required to turn the numerical bounds and certificates into rigorous ones, which (in addition to being additional work) puts additional constraints on the quality of the numerical solution. See, e.g., [33, 17, 23] for examples in the context of SOS certification.
@@ 7.3 Signomial functions and S-cones
Katthän, Naumann and Theobald have shown that, analogously to the SONC results, functions in $\mathcal{S}_{\mathscr{A},\mathscr{B}}$ can be written as sums of AG functions with support $(\mathscr{A},\mathscr{B})$ [18]. They have also characterized both even and odd AG functions in terms of at most two relative entropy cone inequalities [18, Theorem 2.8]. This representation is compatible with the concept of dual certificates. For example, an $n$-variate odd AG function given by (31) with support $(\mathscr{A},\{{\beta}\})$ is nonnegative on $\mathbb{R}^{n}$ if and only if there exists a $\nu\in\mathbb{R}_{+}^{\mathscr{A}}$ such that
@@ 7.4 Nonnegative bases, LP- and SOCP-representable subcones of SOS polynomials
    Polynomials with nonnegative coefficients in a given basis that consists of trivially nonnegative polynomials over $\Delta$, e.g., when each basis polynomial is a product of some of the polynomials $q_{i}$ defining $\Delta$ in (28);
@@ 8 Discussion
We presented a theoretical framework of cone membership certificates that allows us to interpret vectors in the dual cone as certificates of membership in the primal cone. The new certification scheme builds on the classic self-concordance theory of Nesterov and Nemirovski [27], but it goes beyond the well-known result that the gradient map associated with an SCB defines a bijection between the interior of a cone and the interior of its dual. The key difference is that $\mathcal{H}(b)$ and $\mathcal{B}(b)$ are full-dimensional cones, which means that we can use numerical methods to identify any of the certificates, and still establish rigorous certification of membership. Similarly, in the case of a low-dimensional projection $\mathcal{K}_{A}$ of the high-dimensional cone $\mathcal{K}$, an exact preimage $x\in\mathcal{K}$ of a vector $b\in\mathcal{K}_{A}$ can be computed from any (numerically computed) dual certificate. The same infrastructure also allowed us to compute exact primal feasible solutions of general convex conic optimization problems from suitable dual feasible solutions, with duality gaps that approach zero as the dual feasible solutions approach optimality.
We also established that for many cones, some already existing optimization methods can be used to compute dual certificates. Specifically, feasible dual and primal-dual IPAs operating in appropriate neighborhoods of the central path, such as $\mathcal{N}(\eta,\tau)$ defined in (20), automatically work. For certain cones (e.g., linear images of symmetric cones), the existing dual methods automatically compute dual certificates, and these can now be extended to compute exact primal feasible solutions that converge to a primal optimal solution, complementing the dual feasible iterates of the method. For the most general case of nonsymmetric conic optimization, the IPA recently proposed by the authors [29] can be applied.